A disk, initially rotating at 123 rad/s, is slowed down with a constant angular acceleration of magnitude 3.96 rad/s2. (a) How much time does the disk take to stop? (b) Through what angle (rad) does the disk rotate during that time?
Answer #1

initial = 123 rad/sec

= -3.96 rad/sec​​​​2​​​sec2

Using,

= initial + t

0 = 123 - 3.96t

t = 31.06 seconds to stop the disk

(B) using ,

= initial t + 0.5 t2

= (123×31.06) - 0.5 (3.96)(31.06)2

= 3820.38 - 1910.1527

= 1910.22 rad